Find the matrix representation A of the linear function: R2→R3 where

f(x,y)=(5x-y,2x-y,-x+2y) with respect to the standard bases for R2 and R3

1
Expert's answer
2021-08-23T04:49:53-0400

"f(x,y)=(5x-y,2x-y,-x+2y)\\
Canonical\ basis\ of\ \R^3\ is\\
\{(1,0,0),(0,1,0),(0,0,1)\}\\
\begin{bmatrix}
 5 & -1 \\
 2 & -1\\
 -1 & 2
\end{bmatrix}\begin{bmatrix}
 x \\
 y 
\end{bmatrix}=\begin{bmatrix}
 5x-y\\
 2x-y\\
 -x+2y
\end{bmatrix}"